OpenStudy (daniellelovee):

The data in Light-Duty Vehicles (1975–2006) show the efficiency of light-duty vehicles from 1975 (leftmost point of the connected line) to 2006 (rightmost point of the connected line). Each point represents a year in chronological order from the left end to the right.

OpenStudy (daniellelovee):

What generalizations can you make about changes in vehicle efficiency (horsepower multiplied by MPG) over the course of the study?

OpenStudy (daniellelovee):

@Whitemonsterbunny17

OpenStudy (daniellelovee):

@dan815

OpenStudy (daniellelovee):

@mathmate

OpenStudy (mathmate):

Efficiency = Horse Power * MPG Since all coordinates are readable on the graph, you can presumably calculate the efficiency of each model year and draw conclusions from there. Depending on the results, you could even make a straight line or curve model of the efficiency versus year. Another approach would be to plot iso-efficiency curves superimposed on the graph. Each curve would represent an efficiency value, e.g. the curve y=1600/x (1600=13*130) would pass through the point of 1975, and y=4620/x (4620=21*220) would pass through point of 2006, etc. These are inverse variations, the shape of which should be familiar to you.

OpenStudy (daniellelovee):

WOW thank you so much

OpenStudy (mathmate):

You're welcome!

OpenStudy (daniellelovee):

@ichliebedich

OpenStudy (ichliebedich):

1975 and 1982: in 1975, mpg*hp = 1750

OpenStudy (ichliebedich):

1982 it is up to 2200, a 25% increase

OpenStudy (daniellelovee):

yeah but my previous questions were • What generalizations can you make about changes in fuel efficiency (miles per gallon, MPG) over the course of the study? • What generalizations can you make about changes in engine performance (horsepower) over the course of the study? and I thought that with the one I posted I would get the answers to those but I still dont get them :/

OpenStudy (ichliebedich):

do you have answer choices?

OpenStudy (daniellelovee):

i wish lol

OpenStudy (ichliebedich):

i wish you have answer choices too lol

OpenStudy (daniellelovee):

yeah lol do you understand this though?

OpenStudy (ichliebedich):

a little i did have physics but a long time ago

OpenStudy (daniellelovee):

then can you help?

OpenStudy (ichliebedich):

yes just let me figure out the question

OpenStudy (daniellelovee):

alright thank you :)

OpenStudy (ichliebedich):

they are hard questions lol

OpenStudy (daniellelovee):

i knoww lol

OpenStudy (ichliebedich):

Efficiency = Horse Power * MPG, Say your car or truck is rated at 18 mpg. Take the reciprocal of 18 by dividing one by 18 and get 0.055, which is gallons per mile.

OpenStudy (ichliebedich):

horsepower-a unit of power equal to 550 foot-pounds per second (745.7 watts)

OpenStudy (ichliebedich):

i hope that helps :p the questions are a little bit difficult

OpenStudy (daniellelovee):

yeah thank you I hope that is enough for my teacher

OpenStudy (mathmate):

"What generalizations can you make about changes in fuel efficiency (=horse power*MPG) over the course of the study" means what do you see as the general trend of fuel efficiency (defined as Y=HP*MPG) over the years. You already have the graph that shows 1. the year 2. HP and MPG for the year So you need to calculate for a few selected years and tabulate year against Y=HP*MPG. The generalization could be anything like "increases over the years", "decreases over the years", "no significant change", or "irregular variations", or any other description that you noted.